OpenStudy (anonymous):

What is the digit in 10s place in

OpenStudy (anonymous):

\[\Huge 7^{7^{7^{7^{7..1000 Times}}}}\]

OpenStudy (anonymous):

7? lol

OpenStudy (anonymous):

?

OpenStudy (anonymous):

7, because everything is 77777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777777...

OpenStudy (anonymous):

seriously? -.-

OpenStudy (anonymous):

ya lol

OpenStudy (anonymous):

you're wrong with the answer anyway

OpenStudy (anonymous):

lol 8 then?

OpenStudy (anonymous):

nowhere near

OpenStudy (anonymous):

its 7 because it says digit not the entire number

OpenStudy (anonymous):

why are you forcing me to say that yes you are right?

OpenStudy (anonymous):

because i am right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@hartnn bhaiiiiiiiiiiiiii

hartnn (hartnn):

bhai ,relax. aise maths kabhi nai kiya maine .... what i can think of just generalizing.... but thats tedious...idk any definite method for it...

hartnn (hartnn):

if @mukushla has time, he can help...

OpenStudy (♪chibiterasu):

lol i think it's an exponent that you repeat 1000 times.

geerky42 (geerky42):

There is 1000 7s? Juat making sure.

OpenStudy (♪chibiterasu):

If that's the case then it's 7 :I

OpenStudy (anonymous):

i have some idea 7^2=49 7^3=343 7^4=2401 That means always odd..hence 7^7^7..is smt odd therefore 7^odd 49=4k+1 343=4k+3 2401=4k+1 hence 7^4k+3 or 7^4k+1 its 1000 so 7^4k+3 \[\large 7^{4k}7^3\] => 343 x (7^4)^k \[\large 343 (...01)^k\] now 010101 will remain 01 only so we only worry about 343 hence ans=4

geerky42 (geerky42):

I can solve it using brute force by programming...

geerky42 (geerky42):

I will check it. are you sure it is correct?

OpenStudy (anonymous):

yep ans is 4

OpenStudy (♪chibiterasu):

I didn't understand most of that

OpenStudy (anonymous):

@hartnn ??